OpenStudy (anonymous):

What is the slope of the line 5x - 2y = -10?

OpenStudy (compassionate):

Hello, and welcome to OpenStudy! Yo uwant to solve the equation for y, and the middle term, right next to the x, is your slope. y = mx + b, where b is the y-intersect and m is the slope.

OpenStudy (anonymous):

I don't understand it though... Can you help me?

OpenStudy (compassionate):

Subtract 5x and you get -2y = -5x - 10 Now divide by -2

OpenStudy (anonymous):

does y=-5/2? what do I do with the 10

OpenStudy (anonymous):

?

OpenStudy (compassionate):

5x - 2y = -10? -2y = -5x - 10 y = 2/5x + 5

OpenStudy (anonymous):

can you help me with this? Given the equation y − 4 = 3/4(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (compassionate):

Yeah, no problem. You first want to put the equation in slope-intersecpt form. Just solve for y...

OpenStudy (anonymous):

how do you put it into slope-intercept form?

OpenStudy (compassionate):

Solve. For. Y. Just. Add. 4...